# Investigating Conditional Restricted Boltzmann Machines in Regime Detection

### v1

@@ 2.1. Energy-Based Formulation
$2.$ Derivation of Free Energy
@@ 2.3. Dynamic Biases in Conditional RBMs
We define the history vector $v_{<t}$ as the concatenation of the previous NN observations. Given $D$ dimensions per time step, $v_{<t}$ has dimension $N\cdot D$. This effectively conditions the system on a fixed set of past variables using two autoregressive weight matrices: AA (of dimension $ND\times D$) for the visible units, and $B$ (of dimension $ND\times H$) for the hidden units.
@@
In the model this means that for $h_{t}$ we first initialize the hidden node based on the lagged effects on the state and then train on $v_{t}$, effectively setting a prior. This is what theoretically allows, given a suitable NN, to observe volatility clustering effects in the synthetic data.
@@
This formulation allows the CRBM to distinguish between structure that is predictable from the past (captured by AA and $B$) and the innovations or ”surprises” at time $t$ (captured by the interaction weights $W$). Note that in the binary case this probability takes a similar form to the hidden bias inference.
@@ 4.1. Bernoulli-Bernoulli CRBMs
We evaluate the generative fidelity of the Bernoulli-Bernoulli CRBM by examining the Quantile-Quantile ($Q$) plots of real versus synthetic returns. As illustrated in the left panel, the real market data, as expected (Post-2020), exhibits severe leptokurtosis, with extreme quantiles.
@@
Finally, we examine the trajectory of the free energy $F(\mathbf{v})$ over the 2013-2025 period to determine its utility as an unsupervised metric for systemic risk, something first introduced in Section $2.$ Theoretically, a robust CRBM should assign lower energy to stable market states and higher energy to unseen, volatile regimes. However, as illustrated in Figure 3, the free energy signal is dominated by high-frequency noise, rendering it ineffective for regime detection.
@@ 4.2. Gaussian-Bernoulli CRBM Results and Analysis
We evaluate the generative fidelity and the overall market stability implied in the Gaussian-Bernoulli CRBM. The former, we do by examining the $Q$ plots of real vs synthetic returns. As depicted in the figure below, we can clearly observe that the synthetic data collapses data into a gaussian distribution. The model clearly does a good job of fitting the central probability mass but fails at the tails, we posit that this is because the quadratic term in the energy function of the Gaussian-Bernoulli function, in the extreme outliers, effectively forces the $P(\mathbf{v})$ to a standard normal distribution, thus meaning that we, by definition, fail to learn the tails of the distribution.

### v2

@@ 2.1. Energy-Based Formulation
2.2. Derivation of Free Energy
@@ 2.3. Dynamic Biases in Conditional RBMs
We define the history vector $v_{<t}$ as the concatenation of the previous $N$ observations. Given $D$ dimensions per time step, $v_{<t}$ has dimension $N\cdot D$. This effectively conditions the system on a fixed set of past variables using two autoregressive weight matrices: $A$ (of dimension $ND\times D$) for the visible units, and $B$ (of dimension $ND\times H$) for the hidden units.
@@
In the model this means that for $h_{t}$ we first initialize the hidden node based on the lagged effects on the state and then train on $v_{t}$, effectively setting a prior. This is what theoretically allows, given a suitable $N$, to observe volatility clustering effects in the synthetic data.
@@
This formulation allows the CRBM to distinguish between structure that is predictable from the past (captured by $A$ and $B$) and the innovations or ”surprises” at time $t$ (captured by the interaction weights $W$). Note that in the binary case this probability takes a similar form to the hidden bias inference.
@@ 4.1. Bernoulli-Bernoulli CRBMs
We evaluate the generative fidelity of the Bernoulli-Bernoulli CRBM by examining the Quantile-Quantile (QQ) plots of real versus synthetic returns. As illustrated in the left panel, the real market data, as expected (Post-2020), exhibits severe leptokurtosis, with extreme quantiles.
@@
Finally, we examine the trajectory of the free energy $F(\mathbf{v})$ over the 2013-2025 period to determine its utility as an unsupervised metric for systemic risk, something first introduced in Section 2.2. Theoretically, a robust CRBM should assign lower energy to stable market states and higher energy to unseen, volatile regimes. However, as illustrated in Figure 3, the free energy signal is dominated by high-frequency noise, rendering it ineffective for regime detection.
@@ 4.2. Gaussian-Bernoulli CRBM Results and Analysis
We evaluate the generative fidelity and the overall market stability implied in the Gaussian-Bernoulli CRBM. The former, we do by examining the QQ plots of real vs synthetic returns. As depicted in the figure below, we can clearly observe that the synthetic data collapses data into a gaussian distribution. The model clearly does a good job of fitting the central probability mass but fails at the tails, we posit that this is because the quadratic term in the energy function of the Gaussian-Bernoulli function, in the extreme outliers, effectively forces the $P(\mathbf{v})$ to a standard normal distribution, thus meaning that we, by definition, fail to learn the tails of the distribution.